Statistics from a county public health nurse is provided below. Assist her in completing her report by determining the rate for each vital statistic based on the following information.

if a rate can not be calculated, note what information is missing. Provide the rate or information needed in column J, coinciding with the information on that rate. For this county, the midyear population is 500,000, the estimated female population at that time was 75,000.

*Note that all rates are multiplied by 1,000 except Maternal Mortality is by a factor of 10,000.

   Infant deaths: 25
Answer #1

Total mid year population is 500000

Total number of female is 75000

Number of infant death is 25

Missing data is number of live births ,if it was available then we can find out the infant mortality rate

By the given data we can find the crude death rate of infants

Crude death rate = number of deaths /total population ×1000

= 25/ 500000×1000

= 25/500 = 4